OpenStudy (anonymous):

I have a question about Calculus. This problem doesn't seem hard but I'm missing something. s=t^7cost-7tsint-7cost I need to take the derivative

OpenStudy (anonymous):

respect to t?

OpenStudy (anonymous):

= \[7t^6 \cos t- t^7\sin t -7\sin t + 7t \cos t +7 \sin t = \]\[7t^6 \cos t- t^7\sin t + 7t \cos t]